Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
rolland - 16:44 Friday 21 November 2025 (68216) Print this report
Comment to Bad synchronisation of suspension DSPs (68201)

Yesterday afternoon and today, Valerio and the SAT team have reloaded and/or rebooted the different suspension DSPs to re-synchronize them.

As shown in the attached figures, the monitoring channels are back to their nominal values for all the suspensions (including SR).  Note that for online dataDisplay, the channels are available only late in the data collection pipeline, i.e. in the FbmStDy stream, but not in the other online streams. 

For NE and WE marionettes, there is no monitoring and the alert shelving in the DMS ended this week. I have shelved again the 6 associated flags(until January 12th 13h30).

 

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×